LL Incorporated's currently outstanding 11% coupon bonds have a yield to maturity of 13%. LL believes it could issue new bonds at par that would provide a similar yield to maturity. If its marginal tax rate is 30%, what is LL's after-tax cost of debt? Round your answer to two decimal places.

Shi Importers' balance sheet shows $300 million in debt, $50 million in preferred stock, and $250 million in total common equity. Shi's tax rate is 40%, rd = 7%, rps = 6.5%, and rs = 10%. If Shi has a target capital structure of 30% debt, 5% preferred stock, and 65% common stock, what is its WACC? Round your answer to two decimal places.
Answer #1

a.after-tax cost of debt=yield to maturity*(1-tax rate)

=13*(1-0.3)

=9.1%

b.After-tax cost of debt=7*(1-tax rate)

=7*(1-0.4)=4.2%

WACC=Respective costs*Respective weight

=(4.2*0.3)+(6.5*0.05)+(10*0.65)

=8.09%(Approx).
Answer #2

SOLUTION :


LI Company :


After tax cost of debt 

= Pre tax cost of debt * ( 1 - Tax rate in decimals)

= Yield to maturity in percentage * ( 1 - Tax rate in decimals)

= 13 * ( 1 - 0.30)

= 9.10 % (ANSWER)


Shi Importers :


WACC 

= wS * rS + wPS * rPS + wD * rD(1 - T) 

= 0.65 * 10 + 0.05 * 6.5 + 0.30 * 7(1 - 0.40)

= 8.085 = 8.09 % (ANSWER

answered by: Tulsiram Garg